# Quantifying task-relevant representational similarity using decision variable correlation

### v1

@@ 4.1 High trial-by-trial consistency of V4 & IT representations across monkey brains
We combined the neural data from areas V4 and IT, and computed the DVCs. We find that the DVC between the monkeys is about 0.57. We further compute the DVCs for V4 and IT separately, and find the DVC values to be 0.63 and 0.41, respectively. Overall, these results suggest that DVCs across the monkeys’ brain are generally high, implying that the encoding and the decision strategies used by different monkeys are consistent on an image-by-image basis.
We study a set of models (n=43, obtained from Torchvision) [29] pretrained on ImageNet-1k, an influential benchmark in image classification. This also offers a fair comparison between models by controlling for confounding factors related to different training data. We used the same datsets from [28] as above. We feed the images in [28] into deep vision-based neural networks, subject to standard transforms. The model representation is defined as the activation in response to the image, taken from the penultimate layer – the last layer before the final logit layer.
[…]
Network vs. network We next examine the DVCs between different deep neural networks. Specifically, we evaluate DVCs between deep networks from different model families [^3] We define a model family as a set of architectures sharing a canonical computational graph – such as residual, attention, or convolutional block structures—with variation limited to hyperparameters like depth, width, patch size, or token embedding dimension. See Appendix B for more model details. Using DVC, models from the same family or otherwise share architectural similarities are clearly clustered together(Fig. 2a,b). We find that DVCs between models within the same family (similar model structures and training processes) are substantially higher than pairs from different families(p = 1.33e-56, Fig. 2d), aligning with previous findings [30]. We also find DVCs between models not to be exceedingly high. Despite being trained on the same dataset, they do not seem to converge to a single solution, at least not significantly higher compared to the similarity between the two monkeys (Fig. 2d). These results imply that model structures and training processes still play significant roles in the task solutions found by the models.
Notably, these results differ substantially from results obtained by computing error consistency[13, 14]. Geirhos et al. reported that (i) the consistency between model and brain is very low; (ii) the consistency between network models is much higher than the consistency between humans. Later, we will address the difference between the methodologies.
@@ 4.3 Adversarially trained networks, while highly consistent, have low DVCs to the brain
Robustness represents one important difference between deep networks and our perceptual systems. Small perturbations to images that are imperceptible to humans can lead to qualitative errors in deep networks (i.e., adversarial examples) [31]. Adversarial examples reflect the misalignment between representations in deep networks and brains, given certain local perturbations in the inputs. Adversarial robustness can be increased by using adversarial training, e.g., by finding adversarial examples and adding them to the training set. Studies suggest that features learned through adversarial training may be more aligned with human perception [32, 22], posing an intriguing hypothesis that by making networks locally consistent with human perception, network representations may be better aligned with brain representations globally.
@@
These results suggest that adversarially trained models converge to a common solution (despite that these models have different architectures). Their representations diverge from the non-adversarially trained deep networks, and unfortunately, they also diverge from the representation in V4/IT.
@@ 4.4 Networks trained on rich datasets exhibit no increase in DVC to the brain
Whereas ImageNet-1k has been an important benchmark dataset for the image classification community for a decade, recent state-of-the-art models are trained on larger datasets such as ImageNet-21k, which is a scaled-up version of ImageNet-1k, and JFT-300M, which is proprietary. Models trained on larger, more diverse datasets may generalize within a larger domain, and may show better out-of-distribution generalization ability. A recent study showed that models trained on these larger datasets may exhibit better alignment with human behavior [14]. Furthermore, the negative correlations between classification performance and DVCs to the brains (Fig. 2c) suggest the possibility of overfitting to a particular dataset that is much smaller than what brains are trained on evolutionarily, developmentally and during the experiments[34]. Therefore, it is of particular interest to investigate whether models trained on the richer datasets exhibit higher DVCs to the brain.
@@
Figure 5: Comparsion to Cohen’s Kappa. (a) Heatmaps showing the DVC and Cohen’s Kappa for pairs of representative models. (b) There is a strong positive correlation between Cohen’s Kappa and DVC on model-model consistency (evaluated on this dataset) (c) There is a decent positive correlation between Cohen’s Kappa and DVC on model-monkey consistency. (d) The response histogram to ’edges’ distortion based on the model and decision rules used in [13] and the original study [40]. Different colors represent different nerual networks. (e) Simulation results show that Cohen’s Kappa is sensitive to decision biases, while DVC is invariant to decision biases.
@@ 4.5 Comparison to error consistency based on Cohen’s Kappa
One method that is highly relevant to DVC is Cohen’s Kappa. As a classic statistical measure of inter-rater consistency [41], it was recently applied to quantify the error consistency of deep networks and brains [13, 14]. These studies arrived at very different conclusions, namely that model-model similarity is significantly higher than model-human similarity and that models trained on rich datasets are more aligned with humans. At a high level, these results seem to be inconsistent with our findings, because we found that (i) deep networks exhibit modest consistency with the brain; (ii) DVCs between different deep networks trained on ImageNet-1k are not exceedingly large; (iii) models trained on rich datasets have lower DVCs with brains.
@@
What then is causing the difference in the network-network and network-brain consistences between the results reported in  [13] and DVC? According to signal detection theory, Cohen’s Kappa is determined by both the correlation of the underlying decision variables and the decision criterion. Thus, we wondered if the extremely high Cohen’s Kappa values between different networks as reported in  [13] were due the biases in the decisions. We thus performed a second analysis to further investigate the data from  [40] and analysis used in  [13]. Consistent with our hypothesis, we find that the approach in  [13] introduces high decision bias (see Fig.5d) and reduced accuracy, especially when target categories do not align cleanly with the original training labels. We also find that the origin of this large decision bias is because the analysis in [13] is based an aggregated decoder that estimates class probabilities by combining probabilities from related ImageNet-1k classes. Once we substituted the original decoder with a cross-validated logisitic regression classifier, the estimated Cohen’s Kappa values become largely consistent with the DVC we obtained on the main dataset we analyzed. Finally, using a simple simulation, we demonstrated that Cohen’s Kappa is inflated in this setting as bias increases and accuracy drops, whereas DVC remains consistent (Fig.5e).
These results also highlight a key advantage of the DVC method – it is insensitive to the decision biases. The error consistency quantified based on Cohen’s Kappa captures both shared behavior biases and consistency in their underlying decision variables. For future work, it would be interesting to combine the two methods to dissect the contribution of consistency of DVs and the shared biases in the observers. Doing so requires the presence of both neural and behavior data for the same set of stimuli.
@@ 5 Discussion
We have developed a new method, DVC, to quantify the consistency of the two neural representations that emphasizes task-relevant features. DVC is distinct from other popular approaches such as representational similarity analysis [15] or linear regression [5]. Two representations could have high DVC but low consistency according to the representational similarity analysis, or vice versa. For behavioral metrics that aim to characterize the trial-by-trial consistency, one challenge has been to decouple task performance and trial-by-trial consistency. DVC provides a principled way to do so. For future work, it would be interesting to systematically compare and theoretically relate DVC to other proposed similarity measures [15, 5, 18, 17]. We show that DVC reveals surprising negative correlations between (i) the classification performance of the deep networks trained on ImageNet-1k and (ii) the consistency with the neural representation in V4/IT. Also surprisingly, training the deep network adversarially or using rich datasets seem to evoke a decrease, rather than an increase of DVCs. While it is unclear how to close the gap between the image-by-image consistency of the deep networks to that of the brain, we think the following directions might be promising: (i) training networks using datasets that better resemble the stimulus statistics that drives the evolution of the primate visual system [23]; (ii) develop training procedures that better capture the stimulus noise and internal noise of the brain [42], as well as low level properties of the visual system (e.g., optics and foveation).
@@
Limitations First, our results are limited by the number of monkey subjects in the datasets and the number of simultaneously recorded neurons in V4 and IT. Future larger neural datasets would allow for more accurate estimates of DVCs. Second, despite various adjustments in dimensionality reduction and DV decoding that we have experimented with, there may be factors that we have not taken into account that limit the scope and applicability of the results. For example, It is possible that the high-dimensionality of the feature space of some models affected the estimation of the DV. However, the DVCs of these models with the monkeys are not systematically lower, thus it is unlikely that the they are underestimated. Third, while monkeys provide access to neural recordings, the objects shown in the experiments might not have the same behavioral relevance as they do for humans. Thus, caution should be taken when attempting to generalize the result to humans.
@@ A.1 Split normalization recovers true DVC
A.2 Simulation demonstrates the relationship between bias, accuracy and Cohen’s Kappa
In order to demonstrate that this bias in decision could influence Cohen’s Kappa, we did a simple simulation. Suppose there are 10 classes with 100 samples each. The observers output a vector corresponds to the classes. An unbiased perfect observer outputs ’DV (decision variable)’ 1 for the corresponding class and 0 for all other classes (a one-hot vector). For realism, as observers make mistakes, we simply added gaussian noise to the DV output, which results in both lower Cohen’s Kappa and lower DVC. To model a biased imperfect observer, a bias is applied after DV, which is the same for all samples in the same class (e.g. 0.1 for the first class, 0.2 for the second class) etc. Varying bias levels is achieved by scaling the bias added to the output. The final output is one-hot + noise + bias.
Here, Cohen’s Kappa is directly inflated by the shared bias between two observers. On the other hand, because the bias does not affect the underlying DVs, the pre-normalization DVC is unaffected by the addition of bias. However, DVC does become systematically lower when the DVs are dominated by noise. Therefore, Cohen’s Kappa and DVC are distinct in that the former cares about the decision criterion and the latter do not. They can be seen as complimentary in certain scenarios. The simple simulation also hints at the relationship between accuracy, bias and Cohen’s Kappa. We continue this discussion in section D, where we reiterate that Cohen’s Kappa is intimately linked to accuracy.
@@ B.1 Model performances and choices of the penultimate layers
While we do not have a strict criterion on selecting which models to test, we do follow certain principles. First of all, we try to cover a diverse set of model architectures and span the range of model accuracy, which is why we included older models with mediocre performances. Secondly, we try to include models that other studies have previously examined, so it is easier to compare our study to the previous studies. We did exclude some models due to time limits. We intend to examine an even more comprehensive set of models in future work.

### v3

@@ 4.2 Deep networks with higher accuracy on ImageNet exhibit lower DVCs with brains
Network vs. network We next examine the DVCs between different deep neural networks. Specifically, we evaluate DVCs between deep networks from different model families [^3] We define a model family as a set of architectures sharing a canonical computational graph – such as residual, attention, or convolutional block structures—with variation limited to hyperparameters like depth, width, patch size, or token embedding dimension. See Appendix B for more model details. Using DVC, models from the same family or otherwise share architectural similarities are clearly clustered together(Fig. 2a,b). We find that DVCs between pairs of models within the same family (similar model structures and training processes) are substantially higher than pairs from different families (p = 1.33e-56, Fig. 2d), consistent with previous findings [kornblithSimilarityNeuralNetwork2019]. We also find DVCs between models not to be exceedingly high. Despite being trained on the same dataset, they do not seem to converge to a single solution, at least not significantly higher compared to the similarity between the two monkeys (Fig. 2d). These results imply that model structures and training processes still play significant roles in the task solutions found by the models. Notably, these results differ substantially from results obtained by computing error consistency. These studies [geirhosAccuracyQuantifyingTrialbytrial2020, geirhosPartialSuccessClosing2021] reported that (i) the consistency between model and brain is very low; (ii) the consistency between network models is much higher than the consistency between humans. Later, we will address the difference between the methodologies.
@@ 4.3 Adversarially trained networks, while highly consistent, have low DVCs to the brain
Robustness represents one important difference between deep networks and our perceptual systems. Small perturbations to images that are imperceptible to humans can lead to qualitative errors in deep networks (i.e., adversarial examples) [szegedyIntriguingPropertiesNeural2014a]. Adversarial examples reflect the misalignment between representations in deep networks and brains, given certain local perturbations in the inputs. Adversarial robustness can be increased by using adversarial training, e.g., by finding adversarial examples and adding them to the training set. Studies suggest that features learned through adversarial training may be more aligned with human perception [engstromAdversarialRobustnessPrior2019, featherModelMetamersReveal2023], posing an intriguing hypothesis that by making networks locally consistent with human perception, network representations may be better aligned with brain representations globally.
@@ 4.4 Networks trained on rich datasets exhibit no increase in DVC to the brain
Whereas ImageNet-1k has been an important benchmark dataset for the image classification community for a decade, recent state-of-the-art models are trained on larger datasets such as ImageNet-21k, which is a scaled-up version of ImageNet-1k, and JFT-300M, which is proprietary. Models trained on larger, more diverse datasets may generalize within a larger domain, and may show better out-of-distribution generalization ability. A recent study showed that models trained on these larger datasets may exhibit better alignment with human behavior [geirhosPartialSuccessClosing2021]. Furthermore, the negative correlations between classification performance and DVCs to the brains (Fig. 2c) suggest the possibility of overfitting to a particular dataset that is much smaller than what brains are trained on evolutionarily, developmentally, and during the experiments [kaznatcheevNothingMakesSense2022]. Therefore, it is of particular interest to investigate whether models trained on the richer datasets exhibit higher DVCs to the brain.
@@
Figure 5: Comparsion to Cohen’s Kappa. (a) Heatmaps showing the DVC and Cohen’s Kappa for pairs of representative models. (b) There is a strong positive correlation between Cohen’s Kappa and DVC on model-model consistency (evaluated on this dataset) (c) There is a decent positive correlation between Cohen’s Kappa and DVC on model-monkey consistency. (d) The values of Cohen’s Kappa between models (blue) and human subjects (green) are low while Cohen’s Kappa between models and between human subjects are high, consistent with the original report. Re-analyzed based on data from [geirhosImageNettrainedCNNsAre2018]. (e) Scatter plot showing the relationship between Cohen’s Kappa and the difference in accuracy of pairs of observers, and the theoretical upper bound. (f) The response histogram to ’edges’ distortion based on the model and decision rules used in [geirhosAccuracyQuantifyingTrialbytrial2020] and the original study [geirhosImageNettrainedCNNsAre2018]. Different colors represent different nerual networks. (g) Simulation results show that Cohen’s Kappa is sensitive to decision biases, while DVC is invariant to decision biases.
@@ 4.5 Comparison to error consistency based on Cohen’s Kappa
One method that is highly relevant to DVC is Cohen’s Kappa. As a classic statistical measure of inter-rater consistency [cohenCoefficientAgreementNominal1960], Cohen’s Kappa was recently applied to quantify the error consistency of deep networks and brains [geirhosAccuracyQuantifyingTrialbytrial2020, geirhosPartialSuccessClosing2021]. These studies examined human-model alignment using categorical judgements on assorted out-of-distribution stimuli. For example, ‘edges’ where only the edges are illustrated, or ‘silhouettes’ which are filled outlines of objects. These studies arrived at very different conclusions, namely that (i) model-model similarity is significantly higher than human-human similarity, (ii) model-human similarity is extremely low and that (iii) models trained on rich datasets are more aligned with humans. At a high level, these results seem to be inconsistent with our findings, because we found that (i) deep networks exhibit modest consistency with the brain; (ii) DVCs between different deep networks trained on ImageNet-1k are not exceedingly large; (iii) models trained on rich datasets have lower DVCs with brains.
@@
Accuracy difference While Cohen’s Kappa was originally proposed to disentangle the accuracy and consistency of the observers under certain conditions, it remains possible that the two are intermingled in practice [geirhosAccuracyQuantifyingTrialbytrial2020]. To assess this issue, we examine the behavioral responses to ‘edges’ stimuli from  [geirhosImageNettrainedCNNsAre2018]. From the results shown in Fig. 5d,e, it is evident that Cohen’s Kappa is strongly affected by the difference in the accuracy of the two observers [^4] Theoretical upper-bound is given by $\kappa\;\leq\;\frac{(1-d)^{2}}{1+d^{2}}$, where $d$ is the accuracy difference between the two observers. See Appendix A.2 for the derivation.. When the difference in accuracy is high, Cohen’s kappa is low. This suggests that the low error consistency between humans and models reported in  [geirhosAccuracyQuantifyingTrialbytrial2020] is at least partially due to the large differences in accuracy between the two observers. It also potentially explains why models trained on ‘rich’ datasets, which enjoy better out-of-distribution performance, also exhibit better alignment with human behavior based on Cohen’s Kappa [geirhosPartialSuccessClosing2021].
Decoder bias According to signal detection theory, Cohen’s Kappa is determined by both the correlation of the decision variables and decision criterion. Thus, we wondered if the extremely high Cohen’s Kappa values between different networks as reported in [geirhosAccuracyQuantifyingTrialbytrial2020] are due to biases in the decisions. We thus perform a second analysis to further investigate the data from  [geirhosImageNettrainedCNNsAre2018] and analysis used in  [geirhosAccuracyQuantifyingTrialbytrial2020]. Consistent with our hypothesis, we find that the approach in  [geirhosAccuracyQuantifyingTrialbytrial2020] introduces high decision bias (see Fig. 5f) and reduced accuracy, especially when target categories do not align cleanly with the original training labels. We also find that the origin of this large decision bias is because the analysis in [geirhosAccuracyQuantifyingTrialbytrial2020] is based on an aggregated decoder that estimates class probabilities by combining probabilities from related ImageNet-1k classes. Once we substituted the original decoder with a cross-validated logisitic regression classifier, the estimated Cohen’s Kappa values become largely consistent with the DVC we obtained on the main dataset we analyzed. These results suggest that the large Cohen’s Kappa values between different networks are due to the biases in the decoders [^5] See Appendix D.3 for results on the ‘edges’ data using a logistic regression classifier..
These results highlight the advantages of the DVC method. DVC is insensitive to the decision biases, while the error consistency quantified based on Cohen’s Kappa captures both shared behavior biases and consistency in their underlying decision variables. This point is further demonstrated using a simple simulation (see Fig. 5e). Here, we add behavioral biases to the two observers by shifting the decision criteria consistently, so that they both prefer one category over another. Naturally, as the bias increases, the accuracy drops. Cohen’s Kappa strongly reflects these decision biases, whereas DVC remains consistent (Fig.5g). Furthermore, DVC can effectively decouple accuracy v.s. trial-by-trial consistency of the decision variables [sebastianDecisionvariableCorrelation2018].
@@ 5 Discussion
We have developed a method, DVC, to quantify the consistency of two neural representations. The focus on task-relevant features makes DVC different from other popular approaches such as RSA[kriegeskorteRepresentationalSimilarityAnalysis2008a] or linear regression [yaminsUsingGoaldrivenDeep2016]. Two representations may have high DVC yet low consistency according to RSA, or vice versa. For behavioral metrics that aim to characterize trial-by-trial consistency, one challenge has been to decouple task performance and trial-by-trial consistency. DVC provides a principled way to do so. For future work, it would be interesting to combine the analysis of DVC at the behavioral level [sebastianDecisionvariableCorrelation2018] and the neural representation level to dissect the contribution of consistency of DVs and the shared biases of the observers. It would also be interesting to systematically compare DVC to other proposed similarity measures [kriegeskorteRepresentationalSimilarityAnalysis2008a, yaminsUsingGoaldrivenDeep2016, schrimpfBrainScoreWhichArtificial2020, williamsGeneralizedShapeMetrics2022a] [^6] See Appendix E for some preliminary results on comparison to RSA.. Applying the DVC method, we find that there are surprising negative correlations between (i) the classification performance of the deep networks trained on ImageNet-1k and (ii) the consistency with the neural representation in V4/IT. Furthermore, training the deep network adversarially or using rich datasets seems to evoke decrease, rather than increase of DVCs. While it is unclear how to close the gap between the image-by-image consistency of deep networks to that of the brain, the following directions might be promising: (i) training networks using datasets that better resemble the stimulus statistics that drives the evolution of the primate visual system [mehrerEcologicallyMotivatedImage2021]; (ii) develop training procedures that better capture the stimulus noise and internal noise of the brain [dapelloSimulatingPrimaryVisual2020], as well as low level properties of the visual system (e.g., optics and foveation).
@@
Limitations First, our results are limited by the number of monkey subjects in the datasets and the number of neurons recorded simultaneously from V4 and IT. Future larger neural datasets would allow for more accurate estimates of DVCs. Second, despite various adjustments in dimensionality reduction and DV decoding that we have experimented with, there may be factors that we have not taken into account that limit the scope and applicability of the results. For example, It is possible that the high-dimensionality of the feature space of some models affected the estimation of the DV. However, the DVCs of these models with the monkeys are not systematically lower, thus it is unlikely that they are underestimated. Third, while monkeys provide access to neural recordings, the objects shown in the experiments might not have the same behavioral relevance for them as they do for humans. Thus, caution should be taken when attempting to generalize the result to humans.
Acknowledgment We thank Nikolaus Kriegeskorte for helpful discussions. This research was supported by the National Institute Of Neurological Disorders And Stroke of the National Institutes of Health (Award Number R01NS133924) and a Sloan Research Fellowship to X.-X. Wei. The content is solely the responsibility of the authors and does not necessarily represent the official views of the funding agencies.
@@ A.1 Split normalization recovers true DVC
A.2 Upper bound on Cohen’s Kappa with respect to accuracy difference between the two observers
Let $I,J\in\{0,1\}$ be the correctness indicators of two decision makers on the same $N$ items, with accuracies $p_{i}=\mathbb{P}(I=1)$, $p_{j}=\mathbb{P}(J=1)$, and $d=|p_{i}-p_{j}|$. Define
$c_{\mathrm{obs}}=\mathbb{P}(I=J),\qquad c_{\exp}=p_{i}p_{j}+(1-p_{i})(1-p_{j}),\qquad\kappa=\frac{c_{\mathrm{obs}}-c_{\exp}}{1-c_{\exp}}\quad(c_{\exp}<1).$
$c_{\mathrm{obs}}$ is maximized when $I,J$ maximally agree, and $p_{i},p_{j}$ only differ by $d$. Hence $c_{\mathrm{obs}}\leq 1-d$, and
$c_{\mathrm{obs},\max}=1-d.$
For $c_{\exp}$, parametrize the accuracies by their mean $m=\tfrac{p_{i}+p_{j}}{2}$:
$p_{i}=m+\tfrac{d}{2},\qquad p_{j}=m-\tfrac{d}{2},\qquad m\in\big[\tfrac{d}{2},\,1-\tfrac{d}{2}\big].$
Then
$c_{\exp}(m)=p_{i}p_{j}+(1-p_{i})(1-p_{j})=1-(p_{i}+p_{j})+2p_{i}p_{j}=2m^{2}-2m+1-\frac{d^{2}}{2},$
the expression is minimized at $m^{\star}=\tfrac{1}{2}$, yielding
$c_{\exp,\min}=\frac{1}{2}-\frac{d^{2}}{2}.$
Since $\kappa=(c_{\mathrm{obs}}-c_{\exp})/(1-c_{\exp})$ is increasing in $c_{\mathrm{obs}}$ and decreasing in $c_{\exp}$ (for $c_{\exp}<1$), combining the results gives
$\kappa\;\leq\;\frac{(1-d)-\big(\tfrac{1}{2}-\tfrac{d^{2}}{2}\big)}{1-\big(\tfrac{1}{2}-\tfrac{d^{2}}{2}\big)}=\frac{\tfrac{1}{2}-d+\tfrac{d^{2}}{2}}{\tfrac{1}{2}+\tfrac{d^{2}}{2}}=\frac{(1-d)^{2}}{1+d^{2}}.$
Tightness is achieved by $m^{\star}=\tfrac{1}{2}$ (i.e., $p_{i}=\tfrac{1+d}{2}$, $p_{j}=\tfrac{1-d}{2}$) and maximum agreement between $I,J$, which realizes $c_{\mathrm{obs}}=1-d$ and $c_{\exp}=\tfrac{1}{2}-\tfrac{d^{2}}{2}$.
A.3 Simulation demonstrates the relationship between bias, accuracy and Cohen’s Kappa
In order to demonstrate that this decision bias could influence Cohen’s Kappa, we did a simple simulation. Suppose that there are 10 classes with 100 samples each. The observers output a vector which corresponds to the 10 classes. An unbiased perfect observer outputs ’DV (decision variable)’ 1 for the corresponding class and 0 for all other classes (a one-hot vector). For realism, as observers make mistakes, we simply added gaussian noise to the DV output, which results in both lower Cohen’s Kappa and lower DVC. To model a biased imperfect observer, a bias is applied after DV, which is the same for all samples in the same class (e.g. 0.1 for the first class, 0.2 for the second class) etc. Varying bias levels is achieved by scaling the bias added to the output. The final output is one-hot + noise + bias.
Here, Cohen’s Kappa is directly inflated by the shared bias between two observers. On the other hand, because the bias does not affect the underlying DVs, the pre-normalization DVC is unaffected by the addition of bias. However, DVC does become systematically lower when the DVs are dominated by noise. Therefore, Cohen’s Kappa and DVC are distinct in that the former cares about the decision criterion and the latter do not. They can be seen as complementary in certain scenarios. The simple simulation also hints at the relationship between accuracy, bias, and Cohen’s Kappa. We continue this discussion in section D, where we reiterate that Cohen’s Kappa is intimately linked to accuracy.
@@ B.1 Model performances and choices of the penultimate layers
While we do not have a strict criterion for selecting which models to test, we do follow certain principles. First of all, we try to cover a diverse set of model architectures and span the range of model accuracy, which is why we included older models with mediocre performances. Secondly, we try to include models that other studies have previously examined, so it is easier to compare our study with the previous studies. We did exclude some models due to time limits. We intend to examine an even more comprehensive set of models in future work.